OpenStudy (anonymous):

(x-8)(x-k) = x^2 - 5kx +m In the equation above, k and m are constants. If the equation is true for all values of x, what is the value of m?

jimthompson5910 (jim_thompson5910):

First expand out (x-8)(x-k) to get (x-8)(x-k) = x^2 - 5kx + m x^2 - xk - 8x + 8k = x^2 - 5kx + m x^2 - (k + 8)x + 8k = x^2 - 5kx + m Now equate the x terms and solve for k -(k+8)x = -5kx -(k+8) = -5k -k-8 = -5k -8 = -5k+k -8 = -4k -4k = -8 k = -8/(-4) k = 2 Then use this to find m 8k = m 8*2 = m 16 = m m = 16 So the original equation turns into (x-8)(x-k) = x^2 - 5kx + m (x-8)(x-2) = x^2 - 5*2*x + 16 (x-8)(x-2) = x^2 - 10x + 16

OpenStudy (anonymous):

Alright, I almost understand this. My issue comes with "equating the x terms." I'm at this point: -x(k + 8) +8k = -5kx + m How can I just say that -x(k + 8) = -5kx ? And the same with 8k = m ? I'm sure there is a reasonable explanation; but I just don't see it yet.

jimthompson5910 (jim_thompson5910):

the +m is not part of the x terms

jimthompson5910 (jim_thompson5910):

the +m is the constant and that is equal to 8k on the left side

OpenStudy (anonymous):

So just because it's constant it's equal to the left side?

jimthompson5910 (jim_thompson5910):